# Unstable Unlearning: The Hidden Risk of Concept Resurgence in Diffusion Models

### v1

@@ Abstract
Text-to-image diffusion models rely on massive, web-scale datasets. Training them from scratch is computationally expensive, and as a result, developers often prefer to make incremental updates to existing models. These updates often compose fine-tuning steps (to learn new concepts or improve model performance) with “unlearning” steps (to “forget” existing concepts, such as copyrighted works or explicit content). In this work, we demonstrate a critical and previously unknown vulnerability that arises in this paradigm: even under benign, non-adversarial conditions, fine-tuning a text-to-image diffusion model on seemingly unrelated images can cause it to “relearn” concepts that were previously “unlearned.” We comprehensively investigate the causes and scope of this phenomenon, which we term concept resurgence, by performing a series of experiments which compose “mass concept erasure” (the current state of the art for unlearning in text-to-image diffusion models [31]) with subsequent fine-tuning of Stable Diffusion v1.4. Our findings underscore the fragility of composing incremental model updates, and raise serious new concerns about current approaches to ensuring the safety and alignment of text-to-image diffusion models.
@@ 1 Introduction
Modern generative models are not static. In an ideal world, developing new models would require minimal resources, allowing users to tailor unique, freshly trained models to every downstream use case. In practice, making incremental updates to existing models is far more cost-effective, which is why it is common for models developed for one context to be updated for use in another [43, 19, 20]. This paradigm of updating pre-trained models is widely considered beneficial, as it promotes broader and more accessible development of AI. However, for sequential updates to become a sustainable standard, it is critical to ensure that these updates compose in predictable ways. ††⋆Equal Contribution; Corresponding authors: vinithms@mit.edu, ralur@mit.edu
@@
Developers commonly update models to acquire new information or to improve performance—for example, by fine-tuning an existing model on a novel dataset tailored to a particular use case. But developers sometimes seek to remove information from an existing model. One prominent example is machine unlearning, which aims to efficiently update a trained model to “forget” portions of its training data [3, 36, 1], in order to respond to privacy concerns. This is especially important for compliance with regulations like the General Data Protection Regulation (GDPR) “right to be forgotten” [11].
@@
Here, we focus on the related notion of “concept unlearning” in the context of text-to-image diffusion models (hereafter, referred to as “diffusion models”). In contrast to machine unlearning, which targets particular data points, concept unlearning seeks to erase general categories of content, such as offensive or explicit images. There has been substantial recent progress in this area. For example, the current state-of-the-art in “mass concept erasure” (MACE) can now effectively erase dozens of concepts from a pre-trained diffusion model [31]. This is useful in contexts where undesired concepts cannot be comprehensively identified during the pre-training phase, and thus instead must be erased after the model is deployed or as it is adapted for different downstream applications.
@@
Our work begins with a surprising observation: fine-tuning a diffusion model can re-introduce previously erased concepts and this can occur even when fine-tuning is performed on seemingly unrelated concepts (see Figure 1 for a striking yet representative example). This hidden vulnerability, which we call concept resurgence, poses a challenge to the current paradigm of composing model updates via incremental fine-tuning.
In particular, while the current state of the art in concept unlearning may initially suppress the generation of unwanted concepts (e.g., harmful, biased or copyrighted images), a developer cannot presently guarantee that concept unlearning will prevent the accidental reintroduction of these concepts in later updates to the model. As a consequence, consumers who fine-tune a “safe” model might inadvertently reintroduce undesirable content.
@@
This paper systematically explores concept resurgence, identifying it as a critical and previously unrecognized vulnerability in diffusion models. Our primary contributions are as follows:
[…]
Understanding the causes of concept resurgence. We conduct a thorough examination of both the unlearning algorithms and the fine-tuning data involved in concept resurgence. We quantify the extent of the problem across a range of standard benchmarks, finding that the degree of concept resurgence is closely related to the choice of mapping concept (i.e., the more generic / unrelated the concept, the less resurgence) and degree of regularization imposed during unlearning.
Organization of the paper.
Section 2 is dedicated to background and related work. In Section 3, we quantify the extent of concept resurgence across a variety of standard benchmarks. We investigate the influence of unlearning and fine-tuning algorithms in Section 4, followed by an analysis of the role of fine-tuning data in Section 5. Finally, Section 6 discusses limitations and outlines directions for future research.
@@ 2 Background and Related Work
Machine unlearning. We build on a growing literature on machine unlearning [2, 35, 25, 3, 16, 41, 39, 14, 25, 27, 32], which develops methods for efficiently modifying a trained machine learning model to forget some portion of its training data. In the context of classical discriminative models, machine unlearning is often motivated by a desire to preserve the privacy of individuals who may appear in the training data. A key catalyst for this work was the introduction of Article 17 of the European Union General Data Protection Regulation (GDPR), which preserves an individual’s “right to be forgotten” [11]. More recent work in machine unlearning has expanded to include modern generative AI models, which may reproduce copyrighted material, generate offensive or explicit content, or leak sensitive information which appears in their training data [45, 5]. Our work focuses specifically on unlearning in the context of diffusion models [18, 38]. The literature on diffusion models has grown rapidly over the last few years; though we cannot provide a comprehensive overview here, we refer to [45] for an excellent recent survey.
Concept unlearning.
@@
Our work is directly inspired by a line of recent research that proposes methods for inducing models to forget abstract concepts [1, 31, 12, 13], as opposed to simply unlearning specific training examples. A key challenge in this context is maintaining acceptable model performance on concepts that are not targeted for unlearning, especially those closely related to the erased concepts. At the time of this work, Lu et al. [31] (MACE) is the state of the art in terms of both erasure performance and image generation quality after unlearning.
MACE: Mass Concept Erasure in diffusion models.
We build directly on the recent work of Lu et al. [31] for Mass Concept Erasure (MACE) in diffusion models. Broadly speaking, MACE fine-tunes a model to erase certain target phrase (e.g., “an image of a ship”) and their related concepts (e.g., “an image of a boat”) by using a combination of cross-attention refinement and low rank adaptation (LoRA) [20]. Cross-attention refinement modifies the “key” embeddings associated with each token co-existing in the target phrase with the corresponding “key” embedding of co-existing words in a more generic phrase. The second step, LoRA fine-tuning, perturbs the weights of the model to minimize activations in regions which correspond to the target phrase; these regions are identified by segmenting the image using Grounded-SAM [23, 29]. These perturbations are learned via low rank adapatation (LoRA) of the model parameters [20]. Finally, the LoRA modules corresponding to each erased concept are combined to produce a final model by formulating the “fusion” of multiple LoRA modules as a quadratic programming problem. For additional detail on MACE we refer the reader to [31].
Attacking machine unlearning systems.
@@
Finally, a recent line of research explores data poisoning attacks targeting machine unlearning systems, including Chen et al. [7], Marchant et al. [34], Carlini et al. [4], Di et al. [8], Qian et al. [37], Liu et al. [30]. These works show that certain new risks, such as camouflaged data poisoning attacks and backdoor attacks, can be implemented via the “updatability” functionality in machine unlearning, even when the underlying algorithm unlearns perfectly (i.e., simulates retraining-from-scratch). In contrast, our work exposes a qualitatively new kind of vulnerability in machine unlearning, where a previously forgotten concept may be reacquired as a consequence of additional learning.
@@ 3 Composing Updates Causes Concept Resurgence
As discussed in Section 1, the scale of modern diffusion models has motivated a new paradigm in which updates to pretrained models are incrementally composed to avoid retraining models from scratch. These updates broadly take the form of one of two interventions: either the model is updated to learn a new concept, or it is updated to “unlearn” an unwanted concept. The standard procedure for learning new concepts is to curate a dataset of images representing the new concept of interest and fine-tune the model on this dataset. Similarly, to unlearn an unwanted concept(s), an “unlearning” algorithm will typically update the weights of the pretrained model in an attempt to ensure that the model no longer generates content associated with that concept.
These two steps may be repeatedly composed over the lifetime of a deployed model. This paradigm raises an important question:
To what extent is concept erasure robust to compositional updates?
We begin our investigation with Stable Diffusion v1.4, and separately apply MACE to perform a celebrity erasure and object erasure task. We describe these tasks in detail below. For each task, we fine-tune the model on a random set of in-domain concepts after MACE has been applied. For example, in the context of celebrity erasure — where the goal of the erasure task is to “unlearn” the ability to generate images of a particular celebrity — we further fine-tune the resulting model on a random set of celebrity images (which exclude the unlearned celebrity). This is intended to simulate the real world paradigm of composing unlearning with unrelated fine-tuning steps, the latter of which are intended to help the model learn new concepts or otherwise improve performance. In particular, we do not fine-tune the model on adversarially chosen concepts, as our goal is to understand whether benign updates can degrade or otherwise alter performance. For work on adversarial attacks and/or jailbreaking of text-to-image diffusion models, see [33, 44, 9].
@@
Following [31], the first benchmark we consider is inducing the model to forget how to generate certain celebrities (the “erase set”) while retaining the ability to generate others (the “retain set”). We benchmark Stable Diffusion v1.4 on four subtasks in which we apply MACE to unlearn , , or celebrities, and then evaluate whether the model succeeds in generating images of these celebrities (e.g., after being prompted with “A portrait of [erased celebrity name]”). To ensure consistency, both the subtasks and prompts are identical to those in [31]; the full set of celebrities in each subtask, along with the prompts used to evaluate the model, are provided in Section C.1. We quantify model performance across three random seeds by separately computing the mean top-1 accuracy of the Giphy Celebrity Detector (GCD) [17] on both erased and retained celebrities.111The GCD is a popular open source model for classifying celebrity images; [31] document that the GCD achieves top-1 accuracy on celebrity images sampled from Stable Diffusion v1.4. Additionally, for one random seed, we quantify the model’s ability to continue to generate general concepts using the CLIP Score (i.e. the cosine similarity between the prompts and generated images) computed on a random sample of 5000 captions from MSCOCO [28] (COCO-5K) and the FID between the generated images from pretrained Stable Diffusion and those from our experiments for COCO-5K.
Object erasure.

### v2

@@ Abstract
Text-to-image diffusion models rely on massive, web-scale datasets. Training them from scratch is computationally expensive, and as a result, developers often prefer to make incremental updates to existing models. These updates often compose fine-tuning steps (to learn new concepts or improve model performance) with “unlearning" steps (to “forget" existing concepts, such as copyrighted works or explicit content). In this work, we demonstrate a critical and previously unknown vulnerability that arises in this paradigm: even under benign, non-adversarial conditions, fine-tuning a text-to-image diffusion model on seemingly unrelated images can cause it to “relearn" concepts that were previously “unlearned." We comprehensively investigate the causes and scope of this phenomenon, which we term concept resurgence, by performing a series of experiments which compose “concept unlearning" with subsequent fine-tuning of Stable Diffusion v1.4 and Stable Diffusion v2.1. Our findings underscore the fragility of composing incremental model updates, and raise serious new concerns about current approaches to ensuring the safety and alignment of text-to-image diffusion models.
@@ 1 Introduction
Modern generative models are not static. In an ideal world, developing new models would require minimal resources, allowing users to tailor unique, freshly trained models to every downstream use case. In practice, making incremental updates to existing models is far more cost-effective, which is why it is standard for models developed for one context to be updated for use in another [45, 20, 21]. This paradigm of updating pre-trained models is widely considered beneficial, as it promotes broader and more accessible development of AI. However, for sequential updates to become a sustainable standard, it is critical to ensure that these updates compose in predictable ways.
@@
Developers commonly update models to acquire new information or to improve performance—for example, by fine-tuning an existing model on data tailored to a particular use case. But sometimes, developers also seek to remove information from an existing model. One prominent example is machine unlearning, which aims to efficiently update a model to “forget" portions of its training data [4, 32, 2] in order to respond to privacy concerns. This is particularly important to comply with regulations like the General Data Protection Regulation (GDPR) “right to be forgotten" [11].
@@
Here, we focus on the related notion of “concept unlearning" in the context of text-to-image diffusion models (hereafter, referred to as “diffusion models"). In contrast to machine unlearning, which targets individual data points, concept unlearning seeks to erase general categories of content, such as offensive or explicit images. There has been substantial recent progress in this area [14, 27, 16, 13, 48, 22]. For example, the current state-of-the-art algorithms such as “unified concept editing" (UCE) [14] and “mass concept erasure" (MACE) [27] can now effectively erase dozens of concepts from a pre-trained diffusion model. This is useful in contexts where undesired concepts cannot be comprehensively identified during the pre-training phase, and thus instead must be erased after the model is deployed or as it is adapted for different downstream applications.
@@
Our work begins with a surprising observation: fine-tuning a diffusion model can re-introduce previously erased concepts (see Figure 1 for a striking yet representative example). This can occur even when fine-tuning is performed on seemingly unrelated concepts. This hidden vulnerability, which we call concept resurgence, poses a challenge to the current paradigm of composing model updates via incremental fine-tuning. In particular, while the current state of the art in concept unlearning may initially suppress the generation of unwanted concepts (e.g., harmful, biased or copyrighted images), a developer cannot presently guarantee that concept unlearning will prevent the accidental reintroduction of these concepts in later updates to the model. As a consequence, consumers who fine-tune a “safe” model might inadvertently reintroduce undesirable behavior.
@@
This paper systematically explores concept resurgence, identifying it as a critical and previously unrecognized vulnerability in diffusion models. Our primary contributions are:
Demonstrating the prevalence of concept resurgence. Through a series of systematic experiments, we investigate the conditions under which concept resurgence occurs. We show that concept resurgence does not require fine-tuning on data which is similar to the unlearned concept(s), or that the fine-tuning set is chosen adversarially to “jailbreak" the model. Instead, we show that concept resurgence can occur under common and benign usage patterns. Even well-meaning engineers may unintentionally expose users to unsafe or unwanted content that was previously removed. Figure 1 presents a representative example of this phenomenon.
Understanding the severity of concept resurgence. We conduct a thorough examination of different factors that impact the degree of concept resurgence. These include challenges related to scaling unlearning to many simultaneous concepts, and the impact of key implementation choices in common unlearning algorithms.
Investigating the cause(s) of concept resurgence. Finally, we develop a simple toy model to facilitate a systematic investigation into why concept resurgence occurs. This model highlights the fundamental challenges of both detecting and avoiding concept resurgence, and suggests a number of promising avenues for future research.
Organization of the paper. Section 2 covers background and related work. In Section 3, we quantify the extent of concept resurgence across a variety of domains. In Section 4, we explore some of the factors that influence the severity of concept resurgence. Finally, in Section 5 we construct a stylized model to systematically investigate the fundamental drivers of concept resurgence.
@@ 2 Background and related work
Machine unlearning. We build on a growing literature on machine unlearning [3, 31, 24, 4, 17, 43, 39, 15, 24, 25, 28], which develops methods for efficiently modifying a trained machine learning model to forget some portion of its training data. In the context of classical discriminative models, machine unlearning is often motivated by a desire to preserve the privacy of individuals who may appear in the training data. A key catalyst for this work was the introduction of Article 17 of the European Union General Data Protection Regulation (GDPR), which preserves an individual’s “right to be forgotten" [11]. More recent work in machine unlearning has expanded to include modern generative AI models, which may reproduce copyrighted material, generate offensive or explicit content, or leak sensitive information which appears in their training data [47, 6]. Our work focuses specifically on unlearning in the context of text-to-image diffusion models [19, 36]. The literature on diffusion models has grown rapidly over the last few years; though we cannot provide a comprehensive overview here, we refer to [47] for an excellent recent survey.
@@
Concept unlearning. Our work is directly inspired by a line of recent research that proposes methods for inducing models to forget abstract concepts [2, 27, 12, 14, 48, 16, 13, 22], as opposed to simply unlearning specific training examples. A key challenge in this context is maintaining acceptable model performance on concepts that are not targeted for unlearning, especially those closely related to the erased concepts.
We investigate four recently proposed unlearning algorithms: ESD [13], SDD [22], UCE [13], and MACE [27]. At a high level, the first two methods focus on fine-tuning either the cross-attention weights or all of the model parameters such that encountering the concept of interest results in “unconditional" sampling (i.e., sampling which is not conditioned on the unwanted prompt). The latter two used closed-form edits to modify the cross-attention weights – and MACE additionally fine-tunes the remaining model parameters – to erase the concept of interest. We discuss these algorithms in additional detail in Section 4.2.
@@
Attacking machine unlearning systems. Finally, a recent line of research explores data poisoning attacks targeting machine unlearning systems, including [7, 30, 5, 9, 33, 26]. These works show that certain new risks, such as camouflaged data poisoning attacks and backdoor attacks, can be implemented via the “updatability" functionality in machine unlearning, even when the underlying algorithm unlearns perfectly (i.e., simulates retraining-from-scratch). In contrast, our work exposes a qualitatively new kind of vulnerability in machine unlearning, where a previously forgotten concept may be reacquired as a consequence of additional learning.
@@ 3 Composing Updates Causes Concept Resurgence
As discussed in Section 1, the scale of modern diffusion models has motivated a new paradigm in which updates to pretrained models are incrementally composed to avoid retraining models from scratch. These updates broadly take the form of one of two interventions: either the model is updated to learn a new concept, or it is updated to “unlearn" an unwanted concept. The standard procedure for learning new concepts is to curate a dataset of images representing the new concept of interest and fine-tune the model on this dataset. Similarly, to unlearn an unwanted concept(s), an “unlearning" algorithm will typically update the weights of the pretrained model in an attempt to ensure that the model no longer generates content associated with that concept. These two steps may be repeatedly composed over the lifetime of a deployed model. This paradigm raises an important question:
To what extent is concept unlearning robust to compositional updates?
Our investigation into this question begins with four of the most recent and performant unlearning methods discussed in Section 2: MACE, UCE, SDD, and ESD. We apply these unlearning algorithms to four different concept unlearning tasks (celebrity erasure, copyright erasure, unsafe content erasure, and object erasure) and two different diffusion models (Stable Diffusion v1.4 and Stable Diffusion v2.1). We describe these tasks in detail below. For each task, we first apply one of the unlearning algorithms to erase the concept of interest, and then subsequently fine-tune the model on a random set of in-domain concepts. For example, in the context of celebrity erasure — where the goal of the erasure task is to “unlearn" the ability to generate images of a particular celebrity — we further fine-tune the resulting model on a random set of celebrity images (which exclude the unlearned celebrity). This simulates the real world paradigm of composing unlearning with unrelated fine-tuning steps, the latter of which are intended to help the model learn new concepts or otherwise improve performance. In particular, we do not fine-tune the model on adversarially chosen concepts, as our goal is to understand whether benign updates can degrade or otherwise alter performance. For work on adversarial attacks and/or jailbreaking of text-to-image diffusion models, see [29, 46, 10]. Additionally, we focus on settings where the models retained high utility after unlearning.
@@
Celebrity erasure. Following [27], the first benchmark we consider is inducing the model to forget certain celebrities (the “erase set") while retaining the ability to generate others (the “retain set"). We benchmark Stable Diffusion v1.4 and v2.1 in combination with each unlearning algorithm on the task of unlearning celebrities, and then evaluate whether the model succeeds in generating images of these celebrities (e.g., after being prompted with “A portrait of [erased celebrity name]"). To ensure consistency, both the subtasks and prompts are identical to those in [27]; the full set of celebrities in each subtask, along with the prompts used to evaluate the model, are provided in Appendix C. We quantify model performance across three random seeds by separately computing the mean top-1 accuracy of the Giphy Celebrity Detector (GCD) [18] on both erased and retained celebrities.111The GCD is a popular open source model for classifying celebrity images; [27] document that the GCD achieves top-1 accuracy on celebrity images sampled from Stable Diffusion v1.4.
Celebrity | Copyright | |||
Method | Before FT | After FT | Before FT | After FT |
ESD | 0.144 0.011 | 0.950 0.007 | 0.000 0.000 | 0.100 0.067 |
MACE | 0.042 0.004 | 0.391 0.043 | 0.100 0.100 | 0.267 0.167 |
SDD | 0.556 0.203 | 0.965 0.008 | 0.000 0.000 | 0.100 0.033 |
UCE | 0.001 0.001 | 0.004 0.002 | 0.000 0.000 | 0.000 0.000 |
Object | Unsafe | |||
Method | Before FT | After FT | Before FT | After FT |
ESD | 0.192 0.032 | 0.990 0.008 | 0.547 0.073 | 0.840 0.024 |
MACE | 0.045 0.005 | 0.033 0.003 | 0.275 0.058 | 0.319 0.042 |
SDD | 0.000 0.007 | 0.355 0.073 | N/A | N/A |
UCE | 0.023 0.000 | 0.030 0.020 | 0.649 0.010 | 0.670 0.013 |
Copyright erasure. Motivated by recent, well-publicized concerns regarding the ability of diffusion models to generate copyrighted content [40, 41, 44, 49], the second task we consider is one in which we induce the model to unlearn a popular fictional character while retaining the ability to generate other characters. Specifically, we apply each of the four unlearning algorithms to Stable Diffusion v1.4 and v2.1 to unlearn the concept “Iron Man", and then evaluate whether subsequent fine-tuning reintroduces the ability to generate this character (e.g., after being prompted with “a pose of Iron Man in action."). The full set of retained characters and the prompts used to evaluate the model are provided in Appendix C. We quantify the model performance by prompting Molmo 7B-D [8], an open-source multimodal LLM, with the generated image and two questions: “Is [copyrighted character] in this image? Answer Yes or No." and “Who is in this image? State their name only.". We categorize the image as including the character if the response to the first prompt is “Yes" or the character name is correct. We perform this evaluation across three random seeds on the set of evaluation prompts.
Unsafe content erasure. The third task we consider, motivated by concern that diffusion models can generate images containing depictions of self-harm, hate, violence, and/or harassment [37, 35, 34], is the resurgence of unsafe content. We construct this task by leveraging the i2P dataset, which contains a set of prompts that are labeled across different unsafe content categories and their probability of being labeled as inappropriate by the Q16 classifier [38]. As in the previous tasks, we first induce the model to forget the concepts of self-harm, hate, violence, and harassment. We then evaluate whether the model retains the ability to generate these concepts by providing it prompts from the i2P dataset which are labeled as generating an inappropriate image from the unwanted category with a probability of at least 70%. We use the Q16 classifier to evaluate the percentage of unsafe content generated amongst these prompts across three random seeds.